10. Solve the initial value problem using Laplace transform ( 14 points) y + 4y = 2 sin(3t) with y(0) = 1 and y(0) = 1

10. Solve the initial value problem using Laplace transform ( 14 points) y" + 4y = 2 sin(3t) with y(0) = 1 and y'(0) = 1
Answer #1

Solution 9+ 4y = a sin(at) with yea)= If yo)= 1 apply Laplace bansform {{y+4y} = l{Q sinc31)} L {y} + 4L{y} = 21{sin313 5b= GAS? +6BJ2 + 26 A +54B 6- 52C6A+66) + 244+5613 - GA+66= 0 & 264 +54 B = 6 I solve this solve this B - AanĞ (520g)/63200) -